.A person exerts a force of 65N at 520 to the horizontal floor to push a box 5.0m across a floor at constant speed. Calculate the work done on the box. Explain, in terms of energy, what happens to the work done on the box.


1
Expert's answer
2021-01-01T14:06:12-0500

"W=Fs\cos\alpha=65\cdot5\cdot\cos52°=200(J)"


Since the speed does not change, the sum of all forces is zero. The work done is spent on overcoming the force of friction and providing kinetic energy to the body.